OpenStudy (anonymous):

Last Question, Please Help! Jose invests $3,250 at 6% interest compounded annually. What will be the balance in the account after 3.5 years?

OpenStudy (anonymous):

@nothingwasthesame

OpenStudy (anonymous):

I think its 3,250 x .06

OpenStudy (anonymous):

what about the 3.5 years? Where is that?

OpenStudy (anonymous):

em.

OpenStudy (anonymous):

Haven't done these in a while...

OpenStudy (anonymous):

oh.. I know

OpenStudy (anonymous):

so 3250 x .06 every year

OpenStudy (amistre64):

a balance formula might be useful ....

OpenStudy (anonymous):

find that answer out then times it by three. then find half of 3250 x .06 and that will be your answer

OpenStudy (amistre64):

or exceling it

OpenStudy (anonymous):

yes I will make a formula

OpenStudy (anonymous):

umm that doesn't make sense because it says "compounded annually" wouldn't there be a formula for that somewhere?

OpenStudy (amistre64):

\[B_n=B_ok^n-P\frac{1-k^n}{1-k}\]

OpenStudy (anonymous):

amistre64 will help you I got to go :)

OpenStudy (amistre64):

P payments is 0, leaving us the compounding balance

OpenStudy (amistre64):

k is the compounding periodic rate, in this case (1+r/1)

OpenStudy (amistre64):

B(3.5) = 3250(1+.06)^(3.5)

OpenStudy (amistre64):

i was reading it as a loan making payments to start with .. gotta slow down the reading lol

OpenStudy (anonymous):

I got 3985.2345741 and simplified it to 3985.23, is that correct?

OpenStudy (amistre64):

should be, let me dbl chk with a spreadsheet runthru to be sure

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

it seems fine. i was pondering that last half year but it should be fine.

OpenStudy (amistre64):

B(0) = 3250 B(1) = 3250(1.06) B(2) = 3250(1.06)(1.06) B(3) = 3250(1.06)(1.06)(1.06) but then we add only half the interest to it, .03 for half the year B(3.5) = 3250(1.06)(1.06)(1.06)(1.03)

OpenStudy (anonymous):

ok, Thank You :D

OpenStudy (amistre64):

the formula, and the brute math differ by a few cents ...

OpenStudy (amistre64):

good luck